# Topological and statistical signatures for quantifying ambiguity in semantic search

### v1

@@ Abstract
This works reports evidence for the topological signatures of ambiguity in sentence embeddings that could be leveraged for ranking and/or explanation purposes in the context of vector search and Retrieval Augmented Generation (RAG) systems. We proposed a working definition of ambiguity and designed an experiment where we have broken down a proprietary dataset into collections of chunks of varying size - 3, 5, and 10 lines and used the different collections successively as queries and answers sets. It allowed us to test the signatures of ambiguity with removal of confounding factors. Our results show that proxy ambiguous queries (size 10 queries against size 3 documents) display different distributions of homologies 0 and 1 based features than proxy clear queries (size 5 queries against size 10 documents). We then discuss those results in terms increased manifold complexity and/or approximately discontinuous embedding sub-manifolds. Finally we propose a strategy to leverage those findings as a new scoring strategy of semantic similarities.
Neighbourhoods of ambiguous queries have, topologically speaking, a blowfish-like appearance. Image courtesy: Dall-E-3.
Prédire N’est Pas Expliquer (Thom (1992))
(To Predict is not to Explain)
— René Thom, 1991
@@ 1 Introduction
In the field of natural language processing, semantic matching between sentence embeddings often employs metrics such as Euclidean distance, dot product (inner product), or cosine similarity. However, the effectiveness of these measures can be undermined by the fact that the manifold constructed by sentence embeddings is unlikely to be globally smooth, and it may not be locally smooth either. The nature of the semantic manifold, often hypothesized to be low-dimensional based on empirical evidence and intuition, remains elusive yet critical. Literature contains attempts to smooth manifold representation for words and sentences (Hasan and Curry (2017); Kayal (2021); Chu et al. (2023)) but they can be seen as fine-tuning exercises that are by construction limited to a certain domain. The elusive structure of the semantic manifold and its probable discontinuous nature are particularly problematic for ambiguous queries in question-and-answer processes, which may consist in multi-factual or too vague questions. Recent studies have leveraged Topological Data Analysis (TDA, Carlsson and Vejdemo-Johansson (2021)) to interpret the polysemic nature of words (Jakubowski et al. (2020)), indicating local discontinuities within an otherwise smooth word embedding manifold. Building on the concept of polysemy in words, we promote a working and computational definition of ambiguity in sentences based on our experience with vector search and RAG (Gao et al. (2024)) systems in tackling disambiguation problems.
Polysemy refers to a word having multiple meanings, which occurs when a word belongs to more than one semantic domain. Similarly, we examine the relative ambiguity in sentences within a corpus. A corpus $\mathcal{C}$ comprises $n$ sentences, or text chunks, represented as $\mathcal{C}=\{c_{1},...,c_{n}\}$, with each sentence potentially fitting into one or more of $m$ identified semantic domains $S_{\mathcal{C}}=\{s_{1},...,s_{m}\}$, where $m\leq n$. From here onwards, the notation $\mathcal{C}_{/j}$ means the corpus $\mathcal{C}$ with $c_{j}$ removed. Also for every chunk $c_{j}$ we will associate a subset of semantic domains $S_{c_{j}}$. A sentence $c_{j}$ is more ambiguous or polysemic than $c_{k}$ if $|S_{c_{j}}|>|S_{c_{k}}|$.
We now consider a (typically new) sentence or chunk $c^{\prime}$ with a topic set $S_{c^{\prime}}$ polysemic relative to a corpus $\mathcal{C}$ if it meets any of the following working definitions:
  • It belongs to more semantic domains than the sentence in the corpus with the fewest domains:
    $|S_{c^{\prime}}|>\min_{c}(|S_{c}|)\quad\text{with}\quad S_{c^{\prime}}\subset S_{\mathcal{C}}$
  • It introduces at least one semantic domain not present in other sentences:
    $S_{c^{\prime}}\not\subset S_{\mathcal{C}}\quad\text{and}\quad S_{c^{\prime}}\neq\{\emptyset\}$
  • It does not belong to any recognized semantic domain:
    $S_{c^{\prime}}=\{\emptyset\}$
Notice that $c^{\prime}$ may or may not pertain to the original corpus $\mathcal{C}$. In applications, new chunks $c^{\prime}$ represent unseen incoming queries. Any chunk $c_{j}$ will be associated with a single semantic domain $S_{c_{j}}=\{s_{k}\}$ if the set $S$ has been determined, for example, through clustering. It is important to note that clustering of sentence embeddings has been successfully explored with Local Density Approximation or more recently BERTopic (Grootendorst (2022)) to represent semantic domains (called topics). Therefore it follows that relative ambiguity reflects the fact that incoming queries may be associated with multiple pre-existing topics. This can have a degrading impact on RAG systems that depend on clear semantic context for providing answers with high utility for end users. Our goal is to be able to quantify and observe the degree of ambiguity in a sentence $c$ in relation to the corpus $\mathcal{C}$, so we can mitigate or correct the effects of the additional relative ambiguity when retrieving information or generating new content based on an existing context window.
@@
Intuitive definition: An ambiguous sentence is formally defined as a sentence whose embedding satisfies the conditions of connecting to more topics within its corpus than the minimal connections established by other sentences, or not fitting neatly into any predefined topic of the corpus.
@@
    • Sentence 1: "He craned his neck to see the construction site."
    • Sentence 2: "The crane lifted the heavy beams effortlessly."
    • Sentence 3: "A crane flew over the lake at sunset."
  • topics might include {action, machine, bird}.
  • Sentence 1 connects to {action}, Sentence 2 to {machine}, and Sentence 3 to {bird}.
  • To illustrate polysemy further:
    • "He craned his neck to watch the crane lift near the crane."
  • This sentence employs all three topics, making it a polysemic example where the word "crane" connects to {action, machine, bird} simultaneously.
@@
{mdframed}
  1. 1.
    Does an ambiguous sentence embedding display a distinctive geometric and topological signature within its neighborhood, as observed with polysemic word embeddings?
  2. 2.
    And if so, is the signature dependent on a specific training dataset or embedding model architecture?
To address these questions, we conducted a retrieval experiment where we tested queries that either relate to a single or multiple topics covered in the corpus of answers. We designed topological features using TDA toolbox and evaluated their sensitivity to the presence or absence of multiple topics. Key findings include the exploration of the relative nature of sentence ambiguity, demonstrating how the number of concepts contained in the query and corpus affect the occurrence of local manifold discontinuities. We believe our findings can be used to design strategies to improve retrieval in RAG systems.
@@ 2 Related works
Detecting and mitigating effects of queries ambiguities in RAG pipelines is the object of recent contributions. In particular CRAG (Yan et al. (2024)) explicitly categorise in {"Correct","Ambiguous","Incorrect"} a retrieved chunk of text for a given query using a T5-large fine tuned model and subsequently triggers different actionables. That work reports a whole retrieval augmentation pipeline and is an elegant solution but does not give insights about explainability of incorrect retrievals in terms of ambiguity and is domain dependent as it relies on a fine-tuned language model. SELF-RAG (Asai et al. (2023)) also evaluates the relevance of each retrieved chunks given a specific query but by directly using the Large Language Model at generation step. This model is fine-tuned to generate additional relevance tokens to the output in order to trigger generation or a new retrieval loop. As for CRAG, this works focuses on producing flags to iterate the retrieval process but does not interrogate the notion of ambiguity or relevance.
It is also worth mentioning that research groups also question widely used scoring methods such as cosine similarity as unique metric for semantic similarity (Zhou et al. (2022); Steck et al. (2024),) mostly due to words frequency induced biases in training (Wannasuphoprasit et al. (2023)). These questions call for an exploration of additional metrics to characterize the semantic search results. In investigating queries ambiguity, we distinctively focus on an understanding of domain-specific polysemy using topological and geometric methods.
@@ 3 Methodology
In order to investigate topological signatures of ambiguity in a semantic search context we evaluated several embedding models and we built specific features from the query sentence and answer sentences embedding vectors. We opted for a strategy consisting of choosing document chunks for queries and answers but with different chunk sizes which has the advantage of correlating the amount of topics in queries and answers and therefore limiting potential confounding factors in the observed topological signal. Synthetic queries generation does not guarantee such a degree of control to evidence the meaningful signals which explain why we have not chosen this popular dataset generation.
The generic experimental design shown in figure 1 is the following:
  1. 1.
    Chunk Formation: Divide dataset $\mathcal{D}$ into chunks of varying sizes based on the number of text lines ($L\in\mathcal{I}=[3,5,10]$), creating chunk sets $\{\mathcal{C}_{L}\}_{L\in\mathcal{I}}$ (table 1).
  2. 2.
    Setup Use $\mathcal{C}_{x}$ as the corpus $\mathcal{C}$ and $\mathcal{C}_{y}$ as queries set $\mathcal{Q}$. Generate embeddings for both $\mathcal{Q}$ and $\mathcal{C}$. Index $\mathcal{C}$ and generate topics labels via UMAP and HDBSCAN.
  3. 3.
    Retrieval and TDA: Perform semantic search, retrieve top-k=50 corpus chunks. Filter chunks on the conditions $c_{i}\subset q_{j}$ or $c_{i}\supset q_{j}$ where $c_{i}$ is corpus chunk i token sequence and $q_{j}$ is query j token sequence. Tag queries based on multiplicity of topics they relate to.
  4. 4.
    Evaluation: Derive homology based features and identify ambiguous from unambiguous queries signatures in those features.
Figure 1: Experimental pipeline: Documents in dataset $\mathcal{D}$ are divided into chunks of varying sizes to create a collection of chunk sets $\{\mathcal{C}_{L}\}_{L\in\mathcal{I}}$. One set, $\mathcal{C}_{L}$, represents the corpus $\mathcal{C}$ and another serves as query set $\mathcal{Q}$. The topics of the corpus set are defined and used to evaluate the ambiguity of queries from other sets.
3.1 Datasets and Embedding Models
[…]
Dataset chunk size Nchunks
$\mathcal{D}_{BLK}$ 10 2574
- 5 7192
- 3 22404
Table 1: Overview of Datasets sizes
Embedding Model embedding size
SBERT-multiqa(Reimers and Gurevych (2019); Huggingface (2022)) 768
SBERT-msmarco (Reimers and Gurevych (2019); Huggingface (2021)) -
GTR-T5-XL (Ni et al. (2021)) -
Table 2: Overview of Embedding Models and Datasets
We conducted semantic searches using the cosine similarity function for each model and dataset, and collected the top-k = 50 retrieved answers for each query. We embedded the content of these datasets using three different embedding models, as detailed in Table 2 (and appendix C), then stored each resulting vector embeddings in separate FAISS (Douze et al. (2024)) vector store indexes. These models were selected due to the diversity of corpora used in their pre-training. None of these models were fine-tuned on our datasets for the experiment.
3.2 Computational proxy for topics
Despite the established references for word-level polysemy, equivalent references for sentences are not feasible due to complexity and dependency on text corpora. However, we can approximate a sentence semantic domain by clustering sentences as topics in an embedding space. Below, we outline the steps involved using UMAP for dimensionality reduction and HDBSCAN for clustering. Our approach is identical as the one stated in Grootendorst (2022) and more details justifying this approach cn be found there.
Algorithm 1 Identify topics
1:Input: Embedding vectors set $\{\bm{v_{E}}\}$, HDBSCAN parameters, UMAP parameters
2:Output: Topic labels
3:procedure GenerateTopicVectors( $\{\bm{v_{E}}\}$, HDBSCAN params, UMAP params)
4:     $\{\bm{v_{E}}^{\prime}\}\leftarrow\text{UMAP}(\{\bm{v_{E}}\};n_{\text{neighbors}},\min_{\text{dist}},n_{\text{comp}})$
5:     $\bm{S_{\mathcal{C}}}\leftarrow\text{HDBSCAN}(\{\bm{v_{E}}^{\prime}\};m_{\text{size}},m_{\text{samples}},\text{method},\epsilon)$
6:     return $labels=\{l_{1},l_{2},\ldots,l_{n}\}\quad\text{with}\quad l_{i}\in\bm{S_{\mathcal{C}}}\}$
7:end procedure
For hyperparameters tuning and additional details, see Appendix D. The quality of the clusters obtained through HDBSCAN is assessed using silhouette scores, which help evaluate the separation distance between the resulting clusters.
Advantages of the Selected Methods: UMAP is preferred over PCA and t-SNE due to its robust handling of local structures and scalability, making it well-suited for large datasets. HDBSCAN is chosen for its effectiveness in identifying clusters with varying densities, which is crucial for the nuanced clustering required in semantic domain analysis. For a comprehensive understanding of UMAP’s methodology, refer to the seminal paper McInnes et al. (2020).
3.3 Query neighbourhood
@@
Each neighbour of the query relative distance is represented as a scaling factor $\varepsilon_{i}$ w.r.t the relative distance with the nearest neighbour with $\varepsilon_{0}$ = 0.
@@
We opted for this representation as it reflects the local neighbours density which the ranking index $k$ does not. We can then tune an arbitrary scaling factor $\varepsilon_{1}\leq\varepsilon\leq\varepsilon_{k}$ and observe the evolution of topological features. The value of k is taken sufficiently high to allow probing both local and global neighbourhood.
@@
and applying the persistent homology algorithm, as detailed in Appendix A and referenced in Tauzin et al. (2021). Practically, Wasserstein distance and maximum hole lifetime are chosen to quantify $H_{0}$ and $H_{1}$ respectively.

### v2

# Topological quantification of ambiguity in semantic search

@@ Abstract
We studied how the local topological structure of sentence-embedding neighborhoods encodes semantic ambiguity. Extending ideas that link word-level polysemy to non-trivial persistent homology, we generalized the concept to full sentences and quantified ambiguity of a query in a semantic search process with two persistent homology metrics: the 1-Wasserstein norm of $H_{0}$ and the maximum loop lifetime of $H_{1}$.
We formalized the notion of ambiguity as the relative presence of semantic domains or topics in sentences. We then used this formalism to compute ”ab-initio” simulations that encode datapoints as linear combination of randomly generated single topics vectors in an arbitrary embedding space and demonstrate that ambiguous sentences separate from unambiguous ones in both metrics.
Finally we validated those findings with real-world case by investigating on a fully open corpus comprising Nobel Prize Physics lectures from 1901 to 2024, segmented into contiguous, non-overlapping chunks at two granularity: $\sim\!250$ tokens and $\sim\!750$ tokens. We tested embedding with four publicly available models. Results across all models reproduce simulations and remain stable despite changes in embedding architecture.
We conclude that persistent homology provides a model-agnostic signal of semantic discontinuities, suggesting practical use for ambiguity detection and semantic search recall.
Keywords: Semantic Search, Topological Data Analysis, Ambiguity Quantification
@@ 1 Introduction
In the field of natural language processing, semantic matching between sentence embeddings often employs metrics such as Euclidean distance, dot product or cosine similarity. However, the effectiveness of these measures can be undermined by the fact that the manifold constructed by sentence embeddings is unlikely to be globally smooth, and it may not be locally smooth either. The nature of the semantic manifold, often hypothesized to be low-dimensional based on empirical evidence and intuition, remains elusive yet critical. Literature contains attempts to smooth manifold representation for words and sentences (Hasan and Curry (2017); Kayal (2021); Chu et al. (2023)) but they can be seen as fine-tuning exercises that are by construction limited to a certain domain. The elusive structure of the semantic manifold and its probable discontinuous nature are particularly problematic for queries retrieval tasks, which may consist in multi-factual or too vague (ambiguous) questions. Recent studies have leveraged Topological Data Analysis (TDA, Carlsson and Vejdemo-Johansson (2021)) to interpret the polysemic nature of words (Jakubowski et al. (2020)), indicating local discontinuities within an otherwise smooth word embedding manifold. Building on the concept of polysemy in words, we promote a working and computational definition of ambiguity in sentences based on our experience with vector search and RAG (Gao et al. (2024)) systems in tackling disambiguation problems.
Polysemy refers to a word having multiple meanings, which occurs when a word belongs to more than one semantic domains which are mutually exclusive. In other words, given a unique polysemic word, inferring any associated semantic domain (or topics) is equiprobable. Similarly, we examine the relative ambiguity in sentences within a corpus.
@@
Intuitive definition: An polysemic and therefore ambiguous sentence is formally defined as a sentence whose embedding satisfies the conditions of connecting to less topics within its corpus than the minimal number of connections established by other sentences, or not fitting neatly into any predefined topic of the corpus.
@@
    • Sentence 1: ”He craned his neck to see the construction site.”
    • Sentence 2: ”The crane lifted the heavy beams effortlessly.”
    • Sentence 3: ”A crane flew over the lake at sunset.”
    Topics might include {action, machine, bird}. Sentence 1 connects to {action}, Sentence 2 to {machine}, and Sentence 3 to {bird}.
  • To illustrate polysemy:
    • ”I saw a crane.”
    This sentence is ambiguous as it lacks context to ground the word ”crane” in a particular context.
  • A multi-factual query would be:
    • ”He craned his neck to watch the crane lift near the crane.”
    This sentence employs all three topics, making it a multi-factual example where the word ”crane” connects to {action, machine, bird} simultaneously.
A corpus $\mathcal{C}$ comprises $n$ sentences, or text chunks, represented as $\mathcal{C}=\{c_{1},...,c_{n}\}$, with each sentence potentially fitting into one or more of $m$ identified topics $S_{\mathcal{C}}=\{s_{1},...,s_{m}\}$, where $m\leq n$. Also for every chunk $c_{j}$ we will associate a subset of topics $S_{c_{j}}$.
We now consider a (typically new) sentence or chunk $c^{\prime}$ with a topic set $S_{c^{\prime}}$ polysemic relative to a corpus $\mathcal{C}$ if it meets the following working definition:
$|S_{c^{\prime}}|<\min_{c}(|S_{c}|)\quad\text{with}\quad S_{c^{\prime}}\subset S_{\mathcal{C}}$
It contains less topics than the sentence in the corpus with the fewest number of topics. On the contrary, the sentence is a composition of topics (multi-factual) belonging to $S_{\mathcal{C}}$ if:
$S_{c}\subset S_{c^{\prime}}\quad\forall S_{c}\subset S_{\mathcal{C}}$
Finally it is unrelated to $\mathcal{C}$ if it does not belong to any recognized semantic domain:
$S_{c^{\prime}}\cap S_{\mathcal{C}}=\{\emptyset\}$
Notice that in the latter case $c^{\prime}$ does not pertain to the original corpus $\mathcal{C}$. In applications, new chunks $c^{\prime}$ represent unseen incoming queries and in the vast majority of cases they meet polysemic or multi-factual conditions.
@@
1. Does an ambiguous sentence embedding display a distinctive geometric and topological signature within its neighborhood, as observed with polysemic word embeddings? 2. And if so, does the signature depend on a specific training dataset or embedding model architecture?
It is important to note that clustering of sentence embeddings has been successfully explored with Local Density Approximation or more recently BERTopic (Grootendorst (2022)) to represent topics. Therefore it follows that relative ambiguity reflects the fact that incoming queries may be associated with a subset of pre-existing topics present in the corpus. This can have a degrading impact on the precision of RAG systems that depend on clear semantic context for providing answers with high utility for end users. On the other hand, a multi-factual query might be related to multiple elements in the corpus and retrieval ranking system will fail to address complete retrieval of the information (low recall). Our goal is to be able to quantify and observe the degree of ambiguity in a sentence $c$ in relation to the corpus $\mathcal{C}$, so we can mitigate or correct the effects of the additional relative ambiguity when retrieving information or generating new content based on an existing context window.
To address these questions we translated the formal working definitions above into an ”ab-initio” simulation relying solely on the linear representation hypothesis where topics are represented by a set of orthogonal feature vectors encoded in the embedding space and where abstract sentence construction is a linear combination of those features (details in appendix B). This hypothesis as unique axiom for our simulation is strongly motivated by the works of Anthropic on information encoding description (Elhage et al. (2022)) and recent researches that empirically evidence linearly separable representations of semantic domains in large language models hidden states (Saglam et al. (2026)). We designed topological features using Topological Data Analysis toolbox and evaluated their sensitivity to the presence or absence of multiple topics.
We then conducted a retrieval experiment on an open-access corpus consisting of Nobel Prize Physics lectures delivered from 1901 to 2024. The lectures were tokenized and chunked in two granularities (250-token and 750-token windows). We then tested queries that either relate to a single or multiple topics covered in the corpus of answers.
Key findings include the exploration of the relative nature of sentence ambiguity, demonstrating how the number of concepts contained in the query and corpus affect the occurrence of local manifold discontinuities. We believe our findings reinforce the mechanistic approach of information encoding in transformer based LLMs and that it can be used to design strategies to improve retrieval in RAG systems optimizing the precision-recall trade-off.
@@ 2 Related works
Ambiguity created by similarity-metric limitations propagate into retrieval errors in RAG pipelines. Detecting and mitigating those limitations are objects of extensive researches. Several recent works approach this problem: CRAG (Yan et al. (2024)) explicitly categorize in {”Correct”,”Ambiguous”,”Incorrect”} a retrieved chunk of text for a given query using a T5-large fine tuned model and subsequently triggers different actionables. That work reports a whole retrieval augmentation pipeline and is an elegant solution but does not give any insights about explainability of incorrect retrievals in terms of ambiguity and is domain dependent as it relies on a fine-tuned language model. SELF-RAG (Asai et al. (2023)) also evaluates the relevance of each retrieved chunks given a specific query but by directly using the Large Language Model at generation step. This model is fine-tuned to generate additional relevance tokens to the output in order to trigger generation or a new retrieval loop. As for CRAG, this works focuses on producing flags to iterate the retrieval process but does not interrogate the notion of ambiguity or relevance.
It is also worth mentioning that research groups also question widely used scoring methods such as cosine similarity as unique metric for semantic similarity (Zhou et al. (2022); Steck et al. (2024),) mostly due to words frequency induced biases in training (Wannasuphoprasit et al. (2023)). These questions call for an exploration of additional metrics to characterize the semantic search results. In investigating queries ambiguity, we distinctively focus on an understanding of domain-specific polysemy using topological and geometric methods. Very recent work defined geometric based metric: ”the semantic grounding index” by computing angular distance between embeddings with the objective to detect whether large language model answers were similar to the context provided (likely valid answer) or just to the query (likely hallucination) in a RAG framework (Marín (2025)).
We also link our work to the LLM explainability research domain . In particular we are interested in framing the problem in terms of linear representation hypothesis that consists in modeling information features as quasi-orthogonal directions in the model hidden states dimensions. Anthropic published empirical results converging to such a description of information encoding (Elhage et al. (2022)). They also showed, using sparse auto encoders (SAE) (Bricken et al. (2023),Templeton et al. (2024))in their models that such features can be represented by low dimensional manifolds and used to steer generative LLMs outputs in a controlled fashion.Those works have opened the door to research on the features manifold geometries (Engels et al. (2025)],Alexander Modell (2025)).
On a practical aspect, current algorithms focus on ranking retrieved documents to optimize precision and approximate recall through metrics such as Mean Reciprocal Ranking for a predefined top-k documents (MMR@$k$) (Robertson and Zaragoza (2009); Formal et al. (2021); Santhanam et al. (2021)). Little to no attention is given to optimize the true recall of all meaningful datapoints (ie: inferring $k$) and the problem is delegated either:
  • To the documents chunking strategy which can be difficult to validate and is dataset or domain dependent. Moreover chunking cannot account for long range correlation in a document.
  • To an arbitrary cut-off in the ranked results which cumulates dataset dependency and query dependency.
We intend to demonstrate that our working definition of ambiguity is a measurable feature of language and that, in the context of LLM representation of language, it that can compensate the shortcomings of similarity methods.
@@ 3 Methodology
3.1 Persistent homology as a metric.
Persistent homology provides a way to transform the geometric structure of a point cloud into quantitative features. Given a set of points, we gradually increase a connectivity radius and track how the topology of the neighborhood evolves.
At small radii, all points are isolated. As the radius increases, nearby points connect and form clusters. Further increasing the radius may cause clusters to merge or create loop-like structures before eventually collapsing into a single connected component.
The zeroth homology group, $H_{0}$, records the birth and merging of connected components. Its persistence values therefore encode how connectivity emerges across scales and provide a multiscale measure of neighborhood fragmentation.
The first homology group, $H_{1}$, records the appearance and disappearance of loops. Persistent loops indicate the presence of distinct substructures separated by geometric gaps within the neighborhood.
By summarizing these persistence values (through Wasserstein norms for $H_{0}$ and maximum lifetimes for $H_{1}$), we obtain numerical descriptors that reflect the multiscale organization of the query neighborhood. These descriptors serve as topology-aware metrics that complement standard similarity-based measures.
3.2 Evaluating query neighborhood
@@
Each neighbor of the query relative distance is represented as a scaling factor $\varepsilon_{i}$ w.r.t the relative distance with the nearest neighbor with $\varepsilon_{0}$ = 0.
@@
We opted for this representation as it reflects the local neighbors density which the ranking index $k$ does not. We can then tune an arbitrary scaling factor $\varepsilon_{1}\leq\varepsilon\leq\varepsilon_{k}$ and observe the evolution of topological features. The value of k is taken sufficiently high to allow probing both local and global neighborhood.
@@
and applying the persistent homology algorithm, as referenced in Tauzin et al. (2021). Practically, Wasserstein distance and maximum loop lifetime are chosen to quantify $H_{0}$ and $H_{1}$ respectively.